OpenStudy (anonymous):

evaluate the limx->0 sin2x/sin5x with consideration that limx->sinx/x=1

OpenStudy (anonymous):

well for starters, it might help to note that: \[\frac{\sin(2x)}{\sin(5x)} = \frac{\frac{\sin(2x)}{x}}{\frac{\sin(5x)}{x}}\] i dont know if this is the right way to go, just a suggestion.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{\sin2x}{2x}*\frac{5x}{\sin5x}*\frac{2x}{5x}\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}1*1*\frac{2x}{5x}\]\[\frac{2x}{5x}\]

OpenStudy (anonymous):

i know how to do it but my question is that my lecturer says the answer is 2/5 but substitutiting the x in you get zero? so isnt the answer zero or is it 2/5

OpenStudy (anonymous):

but why mathfrick?

OpenStudy (anonymous):

you dont substitute anything in for x, just divide them out: \[\frac{2x}{5x} = \frac{2}{5}(1) = \frac{2}{5}\]

OpenStudy (anonymous):

because the x cancel out okay thank you!...i just saw that lol THANKS SO MUCH JOE!